# cuSZ-$i$: High-Ratio Scientific Lossy Compression on GPUs with Optimized Multi-Level Interpolation

### v4

@@ II Related Work
Recently, many error-bounded lossy compressors for scientific data have been developed, such as SZ [zhao2021optimizing, sz3], ZFP [zfp], QoZ [qoz], SPERR [SPERR], TTHRESH [ballester2019tthresh], and several AI-assisted works [ae-sz, SRN-SZ, han2022coordnet]. Those compressors extend their usability by allowing users to strictly control the accuracy loss in data reconstruction and post-analysis in various domains.
Considering the increasing data generation speed of scientific instruments/simulations, and the popularity of heterogenous HPC systems and applications that directly yield data on GPU platforms, GPU-based scientific lossy compressors have been proposed to deliver orders of magnitude higher throughputs than CPU compressors. Specifically, cuZFP [cuZFP], the CUDA implementation of transform-based ZFP, leverages discrete orthogonal transform and embedding encoding for compression, allowing the user to specify the desired bit rate. Also, prediction-based GPU compressors have been proposed, including cuSZ, cuSZx [szx], FZ-GPU [FZGPU], and cuSZp [cuszp]. While sharing the error-boundness in the reconstructed data, these prediction-based compressors can fit into different use scenarios, as detailed below. 1) cuSZ [cusz2020, cuszplus2021] features fully parallelized prediction-quantization with outlier compacted and coarse-grained Huffman coding that encodes multibyte quantization codes using bits inverse to their frequencies; 2) cuSZx [szx] features a monolithic design to deliver extremely high throughput at the cost of lower data quality and compression ratio. 3) FZ-GPU [FZGPU] alters the compression pipeline by replacing the entire lossless encoding stage with bit-shuffle and dictionary encoding, aiming to deliver higher throughput. 4) cuSZp [cuszp] modifies cuSZ by fusing prediction-quantization and 1D blockwise encoding subroutine into a monolithic GPU kernel, practically achieving high end-to-end throughput. All the high-throughput-oriented GPU-based compressors share the limitation of suboptimal compression ratios. For example, cuSZ, FZ-GPU, and cuSZp all feature Lorenzo prediction, resulting in the ceiled data quality, as described in [zhao2021optimizing]. Furthermore, note that MGARD-X (CUDA-backend) [repo-mgardx] is another GPU compressor practice; though it achieves comparable rate-distortion to cuSZ, it has been reported to be significantly lower in throughput than other GPU-based compressors [FZGPU], making it not fit our use scenarios. Therefore, we exclude it from the discussion.
@@ III Background and Research Motivation
cuSZ [cusz2020, cuszplus2021] is a GPU-based scientific error-bounded lossy compression, which shows state-of-the-art compression ratios and distortions among GPU-based error-bounded scientific lossy compressors. Like the CPU SZ framework, cuSZ also has a modular compression framework composed of data prediction, data quantization, and lossless encoding. In these frameworks, predictions are made for each element of the input data, and the predicted values adjusted by quant-codes (denoted by $q$’s) are accurate with respect to the user-specified error bound and replace the original data values. The quant-codes are either encoded when $|q|<R$, where $R$ is an internal parameter, or otherwise saved as outliers when they are too big (i.e., $|q|\geq R$) for good encoding efficiency. cuSZ features fully parallelized Lorenzo prediction-quantization kernels for compression and decompression. The lossless encoding component features a GPU-based coarse-grained parallelized Huffman encoding, which is the source of the relatively high compression ratios. The critical difference between CPU SZ and cuSZ is that cuSZ does not have a further pass of de-redundancy lossless encoding (e.g., Zstd). This is considered a tradeoff between the throughput and the compression ratio because the sophisticated logic in the redundancy-canceling component can significantly deteriorate the performance of data-processing throughput. We refer readers to the cuSZ papers [cusz2020, cuszp] for more details.
@@ III-B Research Motivations
$\blacksquare$ B.2)  Limitation of existing GPU-based lossy compressors in lossless encoding
Another critical limitation of existing GPU-based lossy compressors is that their lossless encoding modules produce smaller compression ratios due to the concern of compression throughput. Take cuSZ as an example; its lossless encoding module is only the Huffman encoding, mapping each data element to at least 1 bit, and the compression bit rate of cuSZ is always higher than 1. More importantly, the compressed data of cuSZ still has high redundancy in broad cases. Therefore, we need to address this issue by designing or employing new lossless encoding techniques.
@@
§VI-A tweak§VI-B encoding synergycuSZcuSZ-$i$§V spline interpolationw/ outlier filter§V-B G-Interp§V-C auto-tuning§V-D parallelizationhistogrammultibyte Huffmanencodinghigh-speed de-redundancyencoder (e.g., Bitcomp)dual-quant Lorenzo w/ outlier filterhistogrammultibyte Huffmanencodingprediction-quantizationstatisticslossless encoding (1)lossless encoding (2)
Figure 1: The compression pipeline of cuSZ-$i$ and the comparison between it and cuSZ one. The shading indicates the key differences between cuSZ-$i$ (this work) and cuSZ.
@@
  • To maximize the compression ratio boost of cuSZ-$i$ with minor speed degradation, we can enable another lossless encoding pass (e.g., NVIDIA Bitcomp-Lossless) after Huffman encoding.
@@ V-A Basic Design Concept of G-Interp
Fig. 2 demonstrates the basic design of G-Interp with a 3D example. Though accurate, the existing CPU-based interpolation data predictors [zhao2021optimizing, qoz, hpez] are slow (e.g., 0.23 GB/s) and cannot directly be ported to GPU platforms. G-Interp becomes a re-design according to the hardware trait to maximize its data-processing parallelism. Compared to the CPU-based interpolation, it exhibits an accuracy-parallelism tradeoff but still preserves substantially higher data prediction accuracy than the baseline Lorenzo. To these ends, we partition the input data into relatively small chunks and eliminate the data dependencies across them. In the illustration, the 3D input data is partitioned into $8\times 8\times 8$ chunks (or $16\times 16$ for 2D chunks/512 for 1D chunks). Next, to avoid cross-chunk data dependencies, the interpolation operations need to be confined to limited units of data chunks within a short range. Inspired by QoZ [qoz], we introduced anchor points in G-Interp that are losslessly stored so that all interpolations are bounded in the range of any 2 adjacent anchor points (i.e., within one single data chunk). For example, in the 3D case shown in Fig. 2- §VI-A tweak§VI-B encoding synergy1, in each data chunk, one vertex point is assigned as an anchor point (red points in Fig. 2), and 7 other anchor points (on the rest vertices) are borrowed from the surrounding chunks. Afterward, the interpolation within each data chunk is parallelizable and independent of others. In the 3D input data array, approximately 1 of 512 elements becomes anchor points to preserve, whose saving overhead can be further decreased by the additional pass of lossless encoding.
Once we have set up the partitioned data chunks and anchor points, as shown in Fig. 2-( §VI-A tweak§VI-B encoding synergy3§VI-A tweak§VI-B encoding synergy4), the interpolation-based data predictions in each data chunk are performed in parallel. Like in existing works [zhao2021optimizing, qoz], the interpolations are performed level by level, from a large stride to smaller ones (from 4 to 2 and 1 for G-Interp), and each chunk is expanded from $1\times 1\times 1$ to $2\times 2\times 2$, $4\times 4\times 4$, lastly $8\times 8\times 8$. At each level, interpolations are performed subsequently along each dimension. Moreover, within each data chunk, the interpolations on the same level and along the same dimension are also performed in parallel since they have no dependencies. Later, we will feature essential details of this process, including the interpolation splines, error bounds, and the order of interpolation direction.
Figure 2: Design of G-Interp, the GPU-parallelized interpolation-based data predictor.
@@ V-B Interpolation Configurations of G-Interp
As mentioned, because the predicted values with quant-code adjustment replace the original value, the decompression replays the prediction based on the quant-code, where exactly the same interpolation is conducted. Fig. 4 shows another example of the interpolations on a $9\times 9$ 2D grid with an anchor stride of 8. The predicted data points with different interpolation splines are marked with their corresponding shapes. We list each spline function in G-Interp as follows:
@@
§VI-A tweak§VI-B encoding synergy
A $\circ$ $\circ$ * $\bullet$ * $\circ$ $\circ$ A
@@
A $\circ$ $\circ$ * $\bullet$ * $\circ$ $\circ$ A
dim0dim1
A Anchor points
Points predicted by:
$\bullet$ Linear-spline
$\circ$ Quadratic-spline
* Cubic-spline
@@ V-C Profiling-based Auto-tuning of G-Interp Interpolations
We profile the input data array in 2 steps. First, we compute its value range to acquire both the absolute error bound and the value-range-based relative error bound (the absolute error bound divided by the value range). Second, from the data array, G-Interp uniformly samples a small number of data points (e.g., a $\times 4\times$ 4 ×4 sub-grid for 3D cases). For each sampled point, G-Interp performs two instances of cubic spline interpolations along each dimension (e.g., totaling $2\times 3=6$ tests for 3D cases). The profiling kernel accumulates the interpolation prediction errors separately for each interpolation spline and interpolation dimension.
@@ V-D Parallelizing G-Interp on GPU
First, we exploit the data-caching capacity paired with the thread numbers for each thread block. A thread block corresponds to four basic block of $8\times 8\times 8$ elements (featured in §V-A). This builds up a $32\times 8\times 8$ ($x=32$) chunk for a coalescing load (Fig. 2- §VI-A tweak§VI-B encoding synergy2), minimizing the DRAM transaction. After that, borrowed anchor points are loaded from neighbor blocks before interpolation. Fig. 2- §VI-A tweak§VI-B encoding synergy3 shows the first-level interpolation (with a stride of $\alpha_{0}=4$) in the view of a basic block. After the three-stage interpolation in iteration 1 ($\alpha_{0}=4$), the inwardly finer interpolation with stride $\alpha_{1}=2$ continues (Fig. 2- §VI-A tweak§VI-B encoding synergy4) and forth for even finer interpolation with $\alpha_{2}=1$.
Next, each GPU thread block interpolates within a $33\times 9\times 9$ data block, which encloses the $32\times 8\times 8$ input data chunk and borrowed anchor points (Fig. 2- §VI-A tweak§VI-B encoding synergy2). The interpolations are performed as described in §V-A, but now more neighbor points can be available as 4 basic blocks can share their data. For each interpolation level, Fig.2- §VI-A tweak§VI-B encoding synergy5 enumerates the interpolated points along each dimension (suppose the interpolations are performed from dim0 with length 9 to dim2 with length 33). In the level 1 interpolation, the number of interpolated points surpasses the thread number limits of each thread block. Thus, we dynamically assign the number of data points to an active thread.
@@ V-E Advantages of G-Interp: A Quantitative Analysis
First, Fig. 5 compares the quantized prediction errors of cuSZ-Lorenzo, G-Interp, and CPU-interpolation SZ3 (as a benchmark) when applied to field pressure of Miranda dataset [miranda]. In the figure, the nonzero quant-codes (i.e., the prediction error is larger than $eb$) are colored according to the amplitude of the quant-code. Under the same error bounds, the G-Interp design results in much less nonzero quant-codes than cuSZ’s Lorenzo predictor and smaller in amplitude, achieving closer outcomes with the CPU-based SZ3.
Next, because G-Interp exhibits significantly better data prediction accuracy than cuSZ-Lorenzo, its data reconstruction fidelity is also substantially improved over the Lorenzo predictor. In Fig. 6, we propose the comparison between decompression PSNR of G-Interp and cuSZ-Lorenzo (under the same error bound, both CPU-based and GPU-based) on 37 snapshots of dataset RTM (sample 1 snapshot from every 100 timesteps). We find that G-Interp is constantly better than GPU-Lorenzo in terms of PSNR under all error bounds, gaining 2.5 $\sim$ 10 dB of PSNR improvements. Moreover, attributed to the anchor point design, the PSNR from G-Interp even outperforms the CPU version of interpolation (implemented in SZ3 [zhao2021optimizing]).
@@ VI Improving cuSZ-$i$ Lossless Modules
In this section, we detail several important optimizations on cuSZ-$i$’s lossless encoding module, which is crucial in lifting the compression ratio of G-Interp-predicted quant-codes.
@@ VI-A Tweaking Existing Huffman Coding
Adapted from the previous practice [cusz2020], the procedures of building a histogram of quant-code for the Huffman codebook and conducting Huffman coding based on the codebook remain in the cuSZ-$i$ pipeline. Thanks to the more accurate prediction from G-Interp, the histogram contains a more centralized quant-code $q$ distribution. This centralization can be modeled as follows. Recall the outlier-thresholding internal parameter $R$ mentioned in §III-A, for each prediction with respect to the user-specified $eb$, a much smaller range, denoted by $|q|<r^{\bullet}<R$, corresponds to an empirical coverage of very few outliers (e.g., $<1\%$); and at the same $eb$, G-Interp results in a much smaller $r^{\bullet}$ than Lorenzo. As reported in [tian2021revisiting], building a Huffman tree on GPU is worthwhile only when the number of histogram entries (i.e., $r^{\bullet}$) is large enough. Therefore, we shift the procedure of building the Huffman codebook to CPU, with roughly an end-to-end time of 200 us. This time is excluded in the benchmarks in §VII-{C.4, C.5}. Admittedly, this CPU component needs to be further optimized, following approaches such as prebuilding Huffman trees [prebuilt-huffman-tree]. On the other hand, a much smaller $r_{k}<r^{\bullet}$ further signifies this centralization and corresponds to our attempt to allocate small thread-private buffers to cache the count of the center top-$k$ quant-codes. This significantly decreases the transaction between threads’ register files and the shared memory buffer. A large $k$ (or $r_{k}$) can incur higher register pressure; thus, for graceful degradation, $k$ can fall back to $1$, which is still useful for highly compressible cases.

### v5

@@ II Related Work
Recently, many error-bounded lossy compressors for scientific data have been developed, such as SZ [zhao2021optimizing, sz3, hybrid-sz], ZFP [zfp], QoZ [qoz], SPERR [SPERR], TTHRESH [ballester2019tthresh], and several AI-assisted works [ae-sz, SRN-SZ, han2022coordnet]. The usability of these compressors lies in allowing users to strictly control accuracy loss in data reconstruction and post-analysis across various domains. these specialized compressors are also crucial for the treatment of scientific data, one important aspect of which is high-dimensional data interpretation. Previous work [tac2022, amric2023, tacplus2024] emphasize the high-dimensional data continuity, necessitating the preservation of such information. To this end, we are set to design an efficient high-dimensional data compressor.
Moreover, considering the increasing data generation speed of scientific instruments/simulations, and the popularity of heterogenous HPC systems and applications that directly yield data on GPU platforms, GPU-based scientific lossy compressors have been proposed to deliver orders of magnitude higher throughputs than CPU compressors. Specifically, cuZFP [cuZFP], the CUDA implementation of transform-based ZFP, leverages discrete orthogonal transform and embedding encoding for compression, allowing the user to specify the desired bit rate. Also, prediction-based GPU compressors have been proposed, including cuSZ, cuSZx [szx], FZ-GPU [FZGPU], and cuSZp [cuszp]. While sharing the error-boundness in the reconstructed data, these prediction-based compressors can fit into different use scenarios, as detailed below. 1) cuSZ [cusz2020, cuszplus2021] features fully parallelized prediction-quantization with outlier compacted and coarse-grained Huffman coding that encodes multibyte quantization codes using bits inverse to their frequencies; 2) cuSZx [szx] features a monolithic design to deliver extremely high throughput at the cost of lower data quality and compression ratio. 3) FZ-GPU [FZGPU] alters the compression pipeline by replacing the entire lossless encoding stage with bit-shuffle and dictionary encoding, aiming to deliver higher throughput. 4) cuSZp [cuszp] modifies cuSZ by fusing prediction-quantization and 1D blockwise encoding subroutine into a monolithic GPU kernel, practically achieving high end-to-end throughput. All the high-throughput-oriented GPU-based compressors share the limitation of suboptimal compression ratios. For example, cuSZ, FZ-GPU, and cuSZp all feature Lorenzo prediction, resulting in the ceiled data quality, as described in [zhao2021optimizing]. Furthermore, note that MGARD-X (CUDA-backend) [repo-mgardx] is another GPU compressor practice; though it achieves comparable rate-distortion to cuSZ, it has been reported to be significantly lower in throughput than other GPU-based compressors [FZGPU], making it not fit our use scenarios. Therefore, we exclude it from the discussion.
@@ III Background and Research Motivation
cuSZ [cusz2020, cuszplus2021] is a GPU-based scientific error-bounded lossy compression, which shows state-of-the-art compression ratios and distortions among GPU-based error-bounded scientific lossy compressors. Like the CPU SZ framework, cuSZ also has a modular compression framework composed of data prediction, data quantization, and lossless encoding. In these frameworks, predictions are made for each element of the input data, and the predicted values adjusted by quant-codes (denoted by $q$’s) are accurate with respect to the user-specified error bound and replace the original data values. The quant-codes are either encoded when $|q|<R$, an internal parameter, or otherwise saved as outliers when they are too big (i.e., $|q|\geq R$) for efficient encoding. cuSZ features fully parallelized Lorenzo prediction-quantization kernels for compression and decompression. The lossless encoding component features a GPU-based coarse-grained parallelized Huffman encoding, which is the source of the relatively high compression ratios. cuSZ crucially differs from CPU SZ as it does not have a further pass of de-redundancy lossless encoding (e.g., Zstd). This is considered a tradeoff between the throughput and the compression ratio because the sophisticated logic in the redundancy-canceling component can significantly deteriorate the data-processing performance. We refer readers to the cuSZ papers [cusz2020, cuszp] for more details.
@@ III-B Research Motivations
$\blacksquare$ B.2)  Limitation of existing GPU-based lossy compressors due to lossless encoding modules
Another critical limitation of existing GPU-based lossy compressors is that their lossless encoding modules produce smaller compression ratios due to the concern of compression throughput. Taking cuSZ as an example, its sole lossless encoding module, Huffman encoding, uses at least 1 bit to map each data element, making the compression bit rate always higher than 1. More importantly, the compressed data of cuSZ still has high redundancy in broad cases. Therefore, we need to address this issue by designing or employing new lossless encoding techniques.
@@
[…]
Figure 1: The compression pipeline of cuSZ-$i$ and the comparison between it and cuSZ one. The shading indicates the differentiators of cuSZ-$i$ from one design basis, cuSZ.
@@
  • To maximally boost the compression ratio with minor speed degradation, we can enable another lossless encoding pass (e.g., Bitcomp-lossless) after Huffman encoding in cuSZ-$i$.
@@ V-A Basic Design Concept of G-Interp
Fig.2 demonstrates the basic design of G-Interp with a 3D example. Though accurate, the existing CPU-based interpolation data predictors [zhao2021optimizing, qoz, hpez] are slow (e.g., 0.23 GB/s) and cannot directly be ported to GPU platforms. G-Interp becomes a re-design according to the hardware trait to maximize its data-processing parallelism. Compared to the CPU-based interpolation, it exhibits an accuracy-parallelism tradeoff but still preserves substantially higher data prediction accuracy than the baseline Lorenzo. To these ends, we partition the input data into relatively small chunks and eliminate the data dependencies across them. In the illustration, the 3D input data is partitioned into $8^{3}$ chunks (or $16^{2}$ for 2D chunks/512 for 1D chunks). Next, to avoid cross-chunk data dependencies, the interpolation operations need to be confined to limited units of data chunks within a short range. Inspired by QoZ [qoz], we introduced anchor points in G-Interp that are losslessly stored so that all interpolations are bounded in the range of any 2 adjacent anchor points (i.e., within one single data chunk). For example, in the 3D case shown in Fig.2- VI-B1, in each data chunk, one vertex point is assigned as an anchor point (red points in Fig.2), and 7 other anchor points (on the rest vertices) are borrowed from the surrounding chunks. Afterward, the interpolation within each data chunk is parallelizable and independent of others. In the 3D input data array, approximately 1 of 512 elements becomes anchor points to preserve, whose saving overhead can be further decreased by the additional pass of lossless encoding.
Once the partitioned data chunks and anchor points are set up, as shown in Fig.2-( VI-B3VI-B4), the interpolation-based data predictions in each chunk are performed in parallel. Like in existing works [zhao2021optimizing, qoz], the interpolations are performed level by level, from a large stride to smaller ones (from 4 to 2 and 1 for G-Interp), and each chunk is expanded from $1^{3}$ to $2^{3}$, $4^{3}$, lastly $8^{3}$. At each level, interpolations are performed subsequently along each dimension. Moreover, within each chunk, the interpolations on the same level and along the same dimension are also performed in parallel without dependencies. Later, we will feature essential details of this process, including the interpolation splines, error bounds, and the order of interpolation direction.
Figure 2: G-Interp, the parallelized interpolation-based data predictor.
@@ V-B Interpolation Configurations of G-Interp
As mentioned, because the predicted values with quant-code adjustment replace the original value, the decompression precisely replays the interpolation-based prediction from the quant-code. Fig.4 demonstrates the interpolation on a $9\times 9$ 2D grid with an anchor stride of 8. The predicted data points with different interpolation splines are marked with their corresponding shapes. We list each spline function in G-Interp as follows:
@@
VI-B
$\circ$ $\circ$ * $\bullet$ * $\circ$ $\circ$
@@
$\circ$ $\circ$ * $\bullet$ * $\circ$ $\circ$
$\bullet$
$\circ$
@@ V-C Profiling-based Auto-tuning of G-Interp Interpolations
We profile the input data in 2 steps. First, we compute its value range to acquire both the absolute and value-range-based relative error bounds (the former divided by the value range equals to the latter). Second, G-Interp uniformly samples a small number of data points from the input (e.g., a $4^{3}$ sub-grid for 3D cases). For each sampled point, G-Interp performs two instances of cubic spline interpolations along each dimension (e.g., totaling $2\times 3$ tests for 3D cases). The profiling kernel accumulates the interpolation prediction errors separately for each interpolation spline and interpolation dimension.
@@ V-D Parallelizing G-Interp on GPU
First, we exploit the data-caching capacity paired with the thread numbers for each thread block. A thread block corresponds to four basic block of $8^{3}$ elements (featured in §V-A). This builds up a $32_{x}\!\times 8_{y}\!\times 8_{z}$ chunk for a coalescing load (Fig.2- VI-B2), minimizing the DRAM transaction. After that, borrowed anchor points are loaded from neighbor blocks before interpolation. Fig.2- VI-B3 shows the first-level interpolation (with a stride of $\alpha_{0}=4$) in the view of a basic block. After the three-stage interpolation in iteration 1 ($\alpha_{0}=4$), the inwardly finer interpolation with stride $\alpha_{1}=2$ continues (Fig.2- VI-B4) and forth for even finer interpolation with $\alpha_{2}=1$.
Next, each GPU thread block interpolates within a $33_{x}\!\times 9_{y}\!\times 9_{z}$ data block, which encloses the $32_{x}\!\times 8_{y}\!\times 8_{z}$ input data chunk and borrowed anchor points (Fig.2- VI-B2). The interpolations are performed as described in §V-A, but now more neighbor points can be available as 4 basic blocks can share their data. For each interpolation level, Fig.2- VI-B5 enumerates the interpolated points along each dimension (suppose the interpolations are performed from dim0 with length 9 to dim2 with length 33). In the level 1 interpolation, the number of interpolated points surpasses the thread number limits of each thread block. Thus, we dynamically assign the number of data points to an active thread.
@@ V-E Advantages of G-Interp: A Quantitative Analysis
First, Fig.5 compares the quantized prediction errors of cuSZ-Lorenzo, G-Interp, and CPU-interpolation SZ3 (as a benchmark) when applied to field pressure of Miranda dataset [miranda]. In the figure, the nonzero quant-codes (i.e., the prediction error is larger than $eb$) are colored according to the amplitude of the quant-code. Under the same error bounds, the G-Interp design results in much less nonzero quant-codes than cuSZ’s Lorenzo predictor and smaller in amplitude, achieving closer outcomes with the CPU-based SZ3. Next, because G-Interp exhibits significantly better data prediction accuracy than cuSZ-Lorenzo, its data reconstruction fidelity is also substantially improved over the Lorenzo predictor. In Fig.6, we propose the comparison between decompression PSNR of G-Interp and cuSZ-Lorenzo (under the same error bound, both CPU-based and GPU-based) on 37 snapshots of dataset RTM (sample 1 snapshot from every 100 timesteps). We find that G-Interp is constantly better than GPU-Lorenzo in terms of PSNR under all error bounds, gaining PSNR improvements of 2.5 to 10 dB. Moreover, attributed to the anchor point design, the PSNR from G-Interp even outperforms the CPU version of interpolation (implemented in SZ3 [zhao2021optimizing]).
@@ VI Improving cuSZ-$i$ Lossless Modules
In this section, we detail important optimizations on cuSZ-$i$’s lossless encoding module, which crucially boosts the compression ratio of G-Interp-predicted quant-codes.
@@ VI-A Tweaking Existing Huffman Coding
Adapted from the previous practice [cusz2020], the procedures of building a histogram of quant-code for the Huffman codebook and conducting Huffman coding based on the codebook remain in the cuSZ-$i$ pipeline. Thanks to the more accurate prediction from G-Interp, the histogram contains a more centralized quant-code $q$ distribution. This centralization can be modeled as follows. Recall the outlier-thresholding internal parameter $R$ mentioned in §III-A, for each prediction with respect to the user-specified $eb$, a much smaller range, denoted by $|q|<r^{\bullet}<R$, corresponds to an empirical coverage of very few outliers (e.g., $<1\%$); and at the same $eb$, G-Interp results in a much smaller $r^{\bullet}$ than Lorenzo. As reported in [tian2021revisiting], building a Huffman tree on GPU is worthwhile only when the number of histogram entries (i.e., $r^{\bullet}$) is large enough. Therefore, we shift the procedure of building the Huffman codebook to CPU, with roughly an end-to-end time of 200 us. This time is excluded in the benchmarks in §VII-{C.4, C.5}. Admittedly, this CPU component needs to be further optimized, following approaches such as prebuilding Huffman trees [prebuilt-huffman-tree]. On the other hand, a much smaller $r_{k}<r^{\bullet}$ further signifies this centralization and corresponds to our attempt to allocate small thread-private buffers to cache the count of the center top-$k$ quant-codes. This significantly decreases the transaction between threads’ register files and the shared memory buffer. A large $k$ (or $r_{k}$) can incur higher register pressure; thus, for graceful degradation, $k$ can fall back to $1$, which is still helpful for highly compressible cases.